Problem 9-25 Direct Materials and Direct Labor Variances; Computations from Incomplete Data

Sharp Company manufactures a product for which the following standards have been set:

Standard Quantity
or Hours
Standard Price
or Rate
Standard
Cost
Direct materials 3 feet $ 5 per foot $ 15
Direct labor ? hours ? per hour ?

During March, the company purchased direct materials at a cost of $57,510, all of which were used in the production of 3,125 units of product. In addition, 5,100 direct labor-hours were worked on the product during the month. The cost of this labor time was $51,000. The following variances have been computed for the month:

Materials quantity variance $ 1,050 U
Labor spending variance $ 3,500

U

Labor efficiency variance $ 950

U

Required:

1. For direct materials:

a. Compute the actual cost per foot of materials for March.

b. Compute the price variance and the spending variance.

2. For direct labor:

a. Compute the standard direct labor rate per hour.

b. Compute the standard hours allowed for the month’s production.

c. Compute the standard hours allowed per unit of product.
Answer #1

1a) Material quantity variance = (Standard qty-actual qty)Standard price

-1050 = (3125*3*5-5X)

5X = 47925

X(actual quantity) = 9585

Actual Cost per foot = 57510/9585 = 6

1b) Material price variance = (5-6)*9585 = 9585 U

Material spending variance = (3125*15-57510) = 10635 U

2a) Labor rate variance = (Standard rate-actual rate)actual hour

-2550 = (X*5100-51000)

-5100X = -48450

X(Standard rate) = 9.50

2b) Labor efficiency variance = (Standard hour-actual hour)Standard rate

-950 = (X*9.5-5100*9.5)

-9.5X = -47500

X(Standard hour) = 5000 Hour

2c) Standard hour per unit = 5000/3125 = 1.60 hour per unit